Prove that a uniformly continuous function from metric space to another metric space takes cauchy to cauchy sequences


1
Expert's answer
2021-07-18T17:21:27-0400

Let "f: X\to Y" be a uniformly continuous function from a metric space "(X,d_X)" to another metric space "(Y,d_Y)." Let "(x_n)_{n=1}^{\infty}" be a Cauchy sequence. Let "\varepsilon" be arbitrary positive real number. Since the function "f" is uniformly continuous, there exists "\delta>0" such that for any "x,x'\in X" the inequality "d_X(x,x')<\delta" implies "d_Y(f(x),f(x'))<\varepsilon." Since "(x_n)_{n=1}^{\infty}" is a Cauchy sequence, for this "\delta" there exists "n\in\N" such that "d_X(x_k,x_m)<\delta" for any "k\ge n, m\ge n." It follows that "d_Y(f(x_k),f(x_m))<\varepsilon" for any "k\ge n, m\ge n." We conclude that for any "\varepsilon>0" there exists "n\in\N" such that "d_Y(f(x_k),f(x_m))<\varepsilon" for any "k\ge n, m\ge n," and hence "(f(x_n))_{n=1}^{\infty}" is a Cauchy sequence as well.